OpenStudy (anonymous):

Yulian works at the zoo feeding the animals. He puts water in the elephant habitat at the beginning of the day. The table shows how much water, in gallons, remains at different times during the day. What is the equation that represents Yulian’s situation? Record any fractional answers in decimal form. (y – 29) =____(x –____) Yulian’s Situation Hours 2 4 6 8 Water remaining 43 36 29 22

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(y – 29) =__-2/7__(x –___6_) i think but not sure i'm bad ith point slope from

OpenStudy (anonymous):

thanks Alexander sorry if i don't coment right away i am doing two subjects

OpenStudy (chunkymonkay):

the answer was actually (y-29)=-3.5(x-6) you were close alexander